OpenStudy (anonymous):

Find the vertex, the line of symmetry the maximum or minimum value of the quadratic function, and graph the function.

OpenStudy (anonymous):

What's the function?

OpenStudy (anonymous):

f(x)=-x^2+5x+36

OpenStudy (anonymous):

axis of symmetry is at x=5/2

OpenStudy (anonymous):

vertex/maximum (5/2,169/4) axis of symmetry x=5/2

OpenStudy (anonymous):

The x- coordinate of the intercepts are x=?

OpenStudy (anonymous):

factored form: -(x-9)(x+4) so x intercepts are (9,0)&(-4,0)